### v1

# Excited State Densities from Time-Dependent Density Functional Response Theory

Abstract
While the variational principle for excited-state energies leads to a route to obtaining excited-state densities from time-dependent density functional theory, relatively little attention has been paid to the quality of the resulting densities in real space obtained with different exchange-correlation functional approximations, nor how non-adiabatic approximations developed for energies of states of double excitation character perform for their densities. Here we derive an expression directly in real space for the excited-state density, that includes the case of non-adiabatic kernels, and consequently is able, for the first time, to yield densities of states of double-excitation character. Under some well-defined simplifications, we compare the performance of the local-density approximation and exact-exchange approximation, which are in a sense at opposite extremes of the fundamental functional approximations, on local and charge-transfer excitations in one-dimensional model systems, and show that the dressed TDDFT approach gives good densities of double-excitations.
keywords:
American Chemical Society,
[…]
1 Introduction
The advent of time-dependent density functional theory (TDDFT) forty years ago 1, 2, 3 has enabled the calculation of electronic spectra, response properties, and general non-perturbative dynamics, for large and complex systems that would otherwise not be possible. This is due to the one-to-one mapping between time-dependent potentials and densities, for a specified initial state, opening the possibility of finding a non-interacting (Kohn-Sham) system that reproduces the same density while evolving in the computationally simpler one-body Schrödinger equation. While the approach would formally yield exact observables of the interacting system, in practice approximations must be made for the exchange-correlation potential. The majority of applications operate in the linear response regime, where one considers a time-dependent perturbation on the system, from whose response excitation energies and transition properties out of the ground-state can be extracted, and, with the approximations in use today, an unprecedented balance between accuracy and efficiency has been achieved. When thinking about the linear response of a ground-state system, it can be understood that properties related to transition densities between the ground and excited states of the true interacting system, $\langle\Psi_{I}|\hat{n}({\bf r})|\Psi_{0}\rangle$, can be directly obtained, but it is not immediately obvious that excited-state densities $\langle\Psi_{I}|\hat{n}({\bf r})|\Psi_{I}\rangle$ can be extracted from the formalism. However, with the recognition that excited states are stationary points of the Hamiltonian, then using time-independent perturbation theory instead of time-dependent perturbation theory offers a route to extracting their densities from energies computed from linear response; equivalently, framing the problem as a variational one 4, 5. (We note that excited state densities can also be obtained from extrema of the purely the ground-state energy functional 6, 7).
In this paper, we work directly in real space instead of in the space of single-excitations that the earlier pioneering work was framed in 4, 5, and going beyond the use of an adiabatic kernel that the earlier work was limited to. The latter extension enables us to obtain densities for states of double-excitation character. We explore the small-matrix approximation limit of our expressions on a series of one-dimensional model systems that cover local excitations, charge-transfer excitations, and double excitations. Motivated by the works8, 9, 10, 11, 12, 13, 14, 15 suggesting that the hybrid exchange-correlation functionals with a large fraction of exact exchange yield accurate excited-state properties, we explore the performance of the excited-state densities in two opposite limits of approximation – in the adiabatic local density approximation (LDA), and adiabatic exact exchange (EXX). We find that the choice of ground-state approximation is important in getting good Kohn-Sham densities in the first place, on top of which TDDFT provides corrections that bring the density close to the exact. These corrections involve a sum over all orbitals, whose accuracy depends on the choice of exchange-correlation functional, and we study the convergence of this sum. Finally, we show that the DSMA and DSPA frequency-dependent approximations provide good approximations to the densities of states of double-excitation character, while the adiabatic approximation fails.
2 Excited-state densities from time-independent perturbation theory
A standard result from time-independent perturbation theory of quantum mechanics is that the first-order change to an excited-state energy is the expectation value of the perturbation evaluated in that excited state. Applied to many-electron systems when this perturbation is in the externally applied one-body potential, this then yields $\delta E_{I}=\int n_{I}({\bf r})\delta v_{\rm ext}({\bf r})d^{3}r$, where $n_{I}({\bf r})=\sum_{\sigma_{1}..\sigma_{N}}\int|\Psi_{I}({\bf r}\sigma_{1},{\bf r}_{2}\sigma_{2}...{\bf r}_{N}\sigma_{N})|^{2}d^{3}r_{2}..d^{3}r_{N}$ is the density of the unperturbed excited state $I$. Thus, knowledge of the excited state energy and how it varies with the potential yields the excited-state density itself:
$n_{I}({\bf r})=\frac{\delta E_{I}}{\delta v_{\rm ext}({\bf r})}\,.$ (1)
Eq. 1 provides the cornerstone for obtaining excited state densities from TDDFT response: while ground-state DFT provides the ground-state energy $E_{0}$, the TDDFT linear response formalism provides excitation energies $\omega_{I}$, so taking the functional derivative with respect to the external potential of the sum of these yields the excited-state density, exact in principle and dependent on functional approximations made in practise, as we shall see shortly. Thus, we write the excited state energy in two parts:
$E_{I}=E_{0}+\omega_{I}$ (2)
[…]
$\Omega(\omega_{I})\mathbf{G}_{I}=\omega_{I}^{2}\mathbf{G}_{I}$ (3)
where the matrix elements of $\Omega(\omega)$ are expressed in the single-excitation basis, $q=i\rightarrow a,\;q^{\prime}=j\rightarrow b$, as
$\Omega_{qq^{\prime}}(\omega_{I})=\nu^{2}_{q}\delta_{qq^{\prime}}+4\sqrt{\nu_{q}\nu_{q^{\prime}}}f_{{\scriptscriptstyle\rm HXC},qq^{\prime}}[n_{0}](\omega_{I})\,.$ (4)
Here $\nu_{q}={\epsilon}_{a}-{\epsilon}_{i},\;\nu_{q^{\prime}}={\epsilon}_{b}-{\epsilon}_{j}$ are the corresponding KS transition frequencies, and $f_{{\scriptscriptstyle\rm HXC},qq^{\prime}}$ is Hartree-exchange-correlation matrix element, defined through
$f_{{\scriptscriptstyle\rm HXC},qq^{\prime}}[n_{0}](\omega_{I})=\iint d^{3}xd^{3}x^{\prime}\Phi_{q}(\mathbf{x})f_{\scriptscriptstyle\rm HXC}[n_{0}]({\bf x},{\bf x}^{\prime},\omega_{I})\Phi_{q^{\prime}}(\mathbf{x}^{\prime})$ (5)
where
$f_{\scriptscriptstyle\rm HXC}[n_{0}]({\bf x},{\bf x}^{\prime},\omega)=\frac{1}{\left|{\bf x}-{\bf x}^{\prime}\right|}+\int d(t-t^{\prime})e^{{\rm i}\omega(t-t^{\prime})}\left.\frac{\delta v_{\scriptscriptstyle\rm XC}({\bf x},t)}{\delta n({\bf x}^{\prime},t^{\prime})}\right|_{n=n_{0}}$ (6)
and $\Phi_{q}(\mathbf{x})=\phi_{i}(\mathbf{x})\phi_{a}(\mathbf{x})$ represents a KS single-excitation transition density.
Evaluating the derivative in Eq. (1) from Eq. 2 we have
$n_{I}({\bf r})=n_{0}({\bf r})+\frac{\delta\omega_{I}}{\delta v_{\rm ext}({\bf r})}$ (7)
where the first term, the ground-state density, comes from the derivative of the ground-state energy. We focus here on the excited state density difference
$\Delta n_{I}(\mathbf{r})\equiv n_{I}({\bf r})-n_{0}({\bf r})=\frac{\delta\omega_{I}}{\delta v_{\mathrm{ext}}({\bf r})}=\frac{1}{2\omega_{I}}\frac{\delta\omega^{2}_{I}}{\delta v_{\mathrm{ext}}(\mathbf{r})},$ (8)
obtaining the derivative from Eq. (3) in the following way. Multiplying by $\mathbf{G}_{I}^{\dagger}$ on the left, and then taking the derivative, yields the left-hand side as
$\frac{\delta}{\delta v_{\text{ext}}({\bf r})}\left(\mathbf{G}_{I}^{\dagger}\Omega(\omega_{I})\mathbf{G}_{I}\right)$ $=\frac{\delta\mathbf{G}_{I}^{\dagger}}{\delta v_{\text{ext}}({\bf r})}\Omega(\omega_{I})\mathbf{G}_{I}+\mathbf{G}_{I}^{\dagger}\Omega(\omega_{I})\frac{\delta\mathbf{G}_{I}}{\delta v_{\text{ext}}({\bf r})}+\mathbf{G}_{I}^{\dagger}\frac{\delta\Omega(\omega_{I})}{\delta v_{\text{ext}}({\bf r})}\mathbf{G}_{I}$ (9)
$=\omega^{2}_{I}\frac{\delta}{\delta v_{\rm ext}({\bf r})}\left(\mathbf{G}_{I}^{\dagger}\mathbf{G}_{I}\right)+\mathbf{G}_{I}^{\dagger}\frac{\delta\Omega(\omega_{I})}{\delta v_{\rm ext}({\bf r})}\mathbf{G}_{I},$
(where we have used the Hermitian property of $\Omega(\omega)$) while the right-hand-side gives
$\frac{\delta}{\delta v_{\rm ext}({\bf r})}\left(\omega_{I}^{2}\mathbf{G}_{I}^{\dagger}\mathbf{G}_{I}\right)=\omega^{2}_{I}\frac{\delta}{\delta v_{\rm ext}(r)}\left(\mathbf{G}_{I}^{\dagger}\mathbf{G}_{I}\right)+\frac{\delta\omega^{2}_{I}}{\delta v_{\rm ext}(r)}\mathbf{G}_{I}^{\dagger}\mathbf{G}_{I}.$ (10)
Equating the two sides gives
$\Delta n_{I}({\bf r})=\frac{1}{2\omega_{I}}\mathbf{G}_{I}^{\dagger}\frac{\delta\Omega(\omega_{I})}{\delta v_{\rm ext}({\bf r})}\left(\mathbf{G}_{I}^{\dagger}\right)^{-1}=\frac{1}{2\omega_{I}}\mathbf{G}_{I}^{-1}\frac{\delta\Omega(\omega_{I})}{\delta v_{\rm ext}({\bf r})}\mathbf{G}_{I}$ (11)
This expression gives us the exact density-difference, once we compute the functional derivative of the TDDFT matrix (see shortly).
While Eq. (11) is independent of the choice of normalization of the response eigenvectors $\mathbf{G}_{I}$ (as it must be from Eq. (8)), it may be useful to consider the particular normalization derived in Ref. 16 which directly relates the eigenvectors to the oscillator strengths via $f_{I}=\frac{2}{3}\omega_{I}|\langle\Psi_{0}|{\bf r}|\Psi_{I}\rangle|^{2}=\frac{2}{3}\left|\sum_{q,q^{\prime}}\mathbf{r}_{q}^{\dagger}S_{q,q^{\prime}}^{-1/2}G_{I,q^{\prime}}\right|^{2}$ 17, where $S$ is the diagonal matrix $S_{qq^{\prime}}=\delta_{qq^{\prime}}\frac{1}{v_{q}}$. Ref. 16 showed that the required normalization condition is
$\mathbf{G}_{I}^{\dagger}\left(\left.\mathbbm{1}-\frac{\partial\Omega(\omega)}{\partial(\omega^{2})}\right|_{\omega=\omega_{I}}\right)\mathbf{G}_{I}=1.$ (12)
With this choice, we have
$\Delta n_{I}({\bf r})=\frac{1}{2\omega_{I}}\frac{\delta\omega^{2}_{I}}{\delta v_{\text{ext}}({\bf r})}=\frac{1}{2\omega_{I}}\left(\mathbbm{1}+\mathbf{G}_{I}^{\dagger}\left.\frac{\partial\Omega(\omega)}{\partial(\omega^{2})}\right|_{\omega=\omega_{I}}\mathbf{G}_{I}\right)^{-1}\mathbf{G}_{I}^{\dagger}\frac{\delta\Omega(\omega_{I})}{\delta v_{\text{ext}}({\bf r})}\mathbf{G}_{I}.$ (13)
We note that the functional derivative $\frac{\delta\Omega(\omega_{I})}{\delta v_{\rm ext}({\bf r})}$ has two contributions evident from Eq. (4): one is from the variation of KS orbitals and eigenvalues $\{\phi_{i}({\bf r}),\epsilon_{i}\}$ that explicitly appear in Eq. (4) and from the variation of the functional-argument $n_{0}$ of the $f_{\scriptscriptstyle\rm XC}$ kernel under small changes of the external potential, and the other is from the possible frequency-dependence of the xc kernel $f_{\scriptscriptstyle\rm XC}(\omega)$ because this is evaluated at the solution point of Eq. (3), $\omega_{I}$, which depends on $v_{\rm ext}({\bf r})$, $\frac{\delta f_{\scriptscriptstyle\rm XC}(\omega_{I})}{\delta v_{\rm ext}({\bf r})}=\left.\frac{\delta f_{\scriptscriptstyle\rm XC}(\omega)}{\delta\omega}\right|_{\omega=\omega_{I}}\frac{\delta\omega_{I}}{\delta v_{\rm ext}({\bf r})}$. Within the commonly-used adiabatic approximation, the TDDFT matrix is frequency-independent and so $\frac{\delta f_{\scriptscriptstyle\rm XC}(\omega)}{\delta\omega}=0$ and $\frac{\partial\Omega}{\partial(\omega^{2})}$ in Eq. 13 is zero. In the general case, frequency-dependence of the exact kernel brings non-linearity to Eq. (3), resulting therefore in more solutions than the size of the KS single-excitation matrix. This is essential to correctly get states corresponding to states of double-excitation character 20, 21, 22, 23, 24, which are not accessible with an adiabatic approximation. While the frequency-dependence of the $f_{\scriptscriptstyle\rm XC}$ kernel also affects the density-difference dependence through Eq. (13), this equation is, however, still linear, as the matrix is evaluated at the solution frequency $\omega_{I}$, yielding a unique and well-defined density-difference in Eq. 8.
A simplification to Eq. (8) arises once we separate the two contributions to the functional derivative of the $f_{\scriptscriptstyle\rm XC}$ kernel that were mentioned above:
$\frac{\delta}{\delta v_{\rm ext}(\mathbf{r})}f_{\scriptscriptstyle\rm HXC}[n_{0}](\mathbf{r},\mathbf{r}^{\prime},\omega_{I}[n_{0}])=\left.\frac{\delta f_{\scriptscriptstyle\rm HXC}[n_{0}](\mathbf{r},\mathbf{r}^{\prime},\omega)}{\delta v_{\rm ext}(\mathbf{r})}\right|_{\omega=\omega_{I}}+\left.\frac{\partial f_{\scriptscriptstyle\rm HXC}(\mathbf{r},\mathbf{r}^{\prime},\omega)}{\partial\omega^{2}}\right|_{\omega=\omega_{I}}\frac{\delta\omega_{I}^{2}}{\delta v_{\rm ext}(\mathbf{r})}$ (14)
where the second term is only non-zero in the case of non-adiabatic functionals. Transferring this separation to the full $\Omega(\omega)$ matrix, and inserting into Eq. (8), we find
$\Delta n_{I}({\bf r})=\frac{1}{2\omega_{I}}\frac{\delta\omega_{I}^{2}}{\delta v_{\rm ext}(\mathbf{r})}=\frac{1}{2\omega_{I}}\mathbf{G}_{I}^{\dagger}\left.\frac{\delta\Omega(\omega)}{\delta v_{\rm ext}({\bf r})}\right|_{\omega=\omega_{I}}\mathbf{G}_{I}.$ (15)
We note that this does not trivially follow from Eq. 11 because in general $G_{I}$ need not be chosen orthonormal; it certainly is not in the case of the normalization chosen for oscillator strengths, Eqs..
Eq. 11, or Eq. 15 define the exact excited-state densities once the terms of the derivative matrix $\frac{\delta\Omega_{qq^{\prime}}(\omega)}{\delta v_{\rm ext}({\bf r})}$ are calculated. When KS states are considered real, one finds:
$\frac{\delta\Omega_{qq^{\prime}}(\omega)}{\delta v_{\text{ext }}(r)}=$ $\int d^{3}x\left\{2\nu_{q}\Delta n^{\scriptscriptstyle\rm KS}_{q}(\mathbf{x})\delta_{qq^{\prime}}+2f_{{\scriptscriptstyle\rm HXC},qq^{\prime}}\left(\omega\right)\left(\sqrt{\frac{\nu_{q^{\prime}}}{\nu_{q}}}\Delta n^{\scriptscriptstyle\rm KS}_{q}(\mathbf{x})+\sqrt{\frac{\nu_{q}}{\nu_{q^{\prime}}}}\Delta n^{\scriptscriptstyle\rm KS}_{q^{\prime}}(\mathbf{x})\right)\right.$ (16)
$+4\sqrt{\nu_{q}\nu_{q^{\prime}}}\left(\sum_{p\neq i}^{\infty}\frac{1}{\epsilon_{i}-\epsilon_{p}}f_{\mathrm{HXC},\,pa,\,jb}\left(\omega\right)\Phi_{ip}(\mathbf{x})+\sum_{p\neq a}^{\infty}\frac{1}{\epsilon_{a}-\epsilon_{p}}f_{\mathrm{HXC},\,ip,\,jb}\left(\omega\right)\Phi_{pa}(\mathbf{x})\right.$
$\left.\left.+\sum_{p\neq j}^{\infty}\frac{1}{\epsilon_{j}-\epsilon_{p}}f_{\mathrm{HXC},\,ia,\,pb}\left(\omega\right)\Phi_{jp}(\mathbf{x})+\sum_{p\neq b}^{\infty}\frac{1}{\epsilon_{b}-\epsilon_{p}}f_{\mathrm{HXC},\,ia,\,jp}\left(\omega\right)\Phi_{pb}(\mathbf{x})\right)\right\}\left(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}\right)^{-1}(\mathbf{x},{\bf r})$
$+4\sqrt{\nu_{q}\nu_{q^{\prime}}}\int d^{3}x\,\tilde{g}_{{\scriptscriptstyle\rm XC},qq^{\prime}}\left(\mathbf{x},\omega\right)\chi(\mathbf{x},{\bf r})\,.$
Section 1 in the Supplementary Material contains the details of the derivation of Eq. 16. Here we have used the following short-hands. First, we note that when $\chi,\chi_{\scriptscriptstyle\rm S},f_{\scriptscriptstyle\rm HXC}$ are written without any frequency-dependence, this indicates they are to be evaluated at the static limit $\omega=0$: $\chi({\bf r},{\bf r}^{\prime})=\frac{\delta n({\bf r})}{\delta v_{\rm ext}({\bf r}^{\prime})}|_{n=n_{0}},\chi_{\scriptscriptstyle\rm S}({\bf r},{\bf r}^{\prime})=\frac{\delta n({\bf r})}{\delta v_{\scriptscriptstyle\rm S}({\bf r}^{\prime})}|_{n=n_{0}},f_{\scriptscriptstyle\rm HXC}=f_{\scriptscriptstyle\rm HXC}[n_{0}]({\bf r},{\bf r}^{\prime})=\frac{\delta v_{\scriptscriptstyle\rm HXC}[n]({\bf r})}{\delta n({\bf r}^{\prime})}|_{n=n_{0}}$. In contrast, $f_{\scriptscriptstyle\rm HXC}(\omega)=f_{\scriptscriptstyle\rm HXC}[n_{0}]({\bf r},{\bf r}^{\prime},\omega)$ is defined as in Eq. (6). The KS approximation to the single-excitation density difference is denoted $\Delta n^{\scriptscriptstyle\rm KS}_{q}({\bf x})=\phi_{a}^{2}({\bf x})-\phi_{i}^{2}({\bf x})$, while $\Phi_{mn}({\bf x})=\phi_{m}({\bf x})\phi_{n}({\bf x})$ is defined in the same way as the the KS transition density earlier was, except now generalized since $\phi_{m},\phi_{n}$ could both be occupied or unoccupied orbitals. Note that $p$ in the sums run over all states, occupied and unoccupied, so this expression probes the xc kernel matrix elements (Eq. 5) in a much larger range than needed in the calculation of the excitation spectra; the latter involves matrix elements only between single excitations. The real-space matrix $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}({\bf x},{\bf r})$ is defined using the static $f_{\scriptscriptstyle\rm HXC}$ kernel and KS response function, and the product $f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}$ indicates the convolution, $f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}({\bf x},{\bf r})=\int d^{3}x^{\prime}f_{\scriptscriptstyle\rm HXC}({\bf x},{\bf x}^{\prime})\chi_{\scriptscriptstyle\rm S}({\bf x}^{\prime},{\bf r})$. Finally, $\tilde{g}_{{\scriptscriptstyle\rm HXC},qq^{\prime}}({\bf x})=\int d^{3}x^{\prime}d^{3}x^{\prime\prime}\tilde{g}_{\scriptscriptstyle\rm XC}[n_{0}](\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime},\mathbf{x},\omega)\Phi_{q}({\bf x}^{\prime})\Phi_{q^{\prime}}({\bf x}^{\prime\prime})$ is the matrix in KS single-excitation space of
$\tilde{g}_{\scriptscriptstyle\rm XC}[n_{0}](\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime},\mathbf{x},\omega)=\left.\frac{\delta f_{\scriptscriptstyle\rm XC}[n_{0}^{\prime}](\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime},\omega)}{\delta n_{0}^{\prime}(\mathbf{x})}\right|_{n_{0}^{\prime}=n_{0}}$ (17)
We note that this is a distinct object from the second-order response kernel that appears in quadratic response theory, which is the Fourier transform of the functional derivative $g_{\scriptscriptstyle\rm XC}[n_{0}](\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime},\mathbf{x},t-t^{\prime},t-t^{\prime\prime})=\left.\frac{\delta f_{\scriptscriptstyle\rm XC}[n_{0}^{\prime}](\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime},t-t^{\prime})}{\delta n_{0}^{\prime}(\mathbf{x},t-t^{\prime\prime})}\right|_{n_{0}^{\prime}=n_{0}}$.
Eqs. 1516, once added to the ground-state density, give in-principle the exact excited-state density for the state $I$. As such Eq.(16) must integrate to zero. To see that it does, consider the Taylor expansions of the inverse
$(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}=\mathbbm{1}+f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}+f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}+\dots$ (18)
and that for the Dyson equation for $\chi({\bf x},{\bf r})$ (see also the Supplementary Material). One then observes that the ${\bf r}$-dependence in all the terms in $\frac{\delta\Omega_{qq^{\prime}}(\omega)}{\delta v_{\text{ext }}({\bf r})}$ involve KS density differences $\Delta n^{\scriptscriptstyle\rm KS}_{q}({\bf r})$ or KS transition densities $\Phi_{mn}({\bf r})$, and thus integrate to zero.
We note that Refs. 25, 26, 4, 5 begin instead with finding stationary points of a Lagrangian that expresses the variational property of the excited state energy, working in the KS single-excitation subspace to directly obtain matrix equations for the so-called $Z$-vector that contains relaxation contributions to a difference density matrix; solving these equations yields excited state gradients and other properties. Our Eqs. 1516 are equivalent but instead yields the densities directly in real-space, and moreover, applies also to the case of non-adiabatic kernels.
2.1 Small-Matrix Approximation
While Eqs. 1516 provide the exact prescription for the excited-state density difference, we will make a diagonal approximation to the TDDFT matrix Eq. (3) to simplify the calculation in our examples. Neglecting the mixing between different single KS excitations is known as the Small-Matrix Approximation (SMA) 19, and taking the diagonal element of Eq. 3 yields
${\omega^{2}_{I}}\;\;{\overset{\rm SMA}{\approx}}\;\;\nu_{q}^{2}+4\nu_{q}{f_{\scriptscriptstyle\rm HXC}}_{qq}(\omega_{I}).$ (19)
Under this approximation, with normalization condition Eq. (12), Eq. (15) reduces to
$\Delta n_{I}^{\mathrm{SMA}}({\bf r})$ $=\frac{G_{I}^{2}}{2\omega_{I}}\left.\frac{\delta\Omega_{qq}(\omega)}{\delta v_{\rm ext}({\bf r})}\right|_{\omega=\omega_{I}}$ (20)
with the eigenvector $G_{I}$ reducing to the number
$G_{I}=\frac{1}{\sqrt{1-\left.\frac{\partial\Omega_{qq}(\omega)}{\partial\omega^{2}}\right|_{\omega=\omega_{\text{SMA},I}}}}$ (21)
Inserting the diagonal elements of Eq. (16) directly into Eq. (20) leads to the SMA density-difference:
$\Delta n_{I}^{\mathrm{SMA}}(\mathbf{r})$ $=\frac{{G_{I}^{2}}}{\omega_{I}}\int d^{3}x\left\{\left[\left(\nu_{q}+2f_{\mathrm{Hxc}qq}(\omega_{I})\right)\Delta n_{I}^{\mathrm{KS}}(\mathbf{x})\right.+4\sum_{p\neq a}^{\infty}\frac{\nu_{q}}{\epsilon_{a}-\epsilon_{p}}{f_{\mathrm{Hxc}\,ip,ia}(\omega_{I})}{\Phi_{pa}\left(\mathbf{x}\right)}\right.$ (22)
$\left.-4\sum_{p\neq i}^{\infty}\frac{\nu_{q}}{\epsilon_{p}-\epsilon_{i}}{f_{\mathrm{Hxc}\,pa,ia}(\omega_{I})}{\Phi_{ip}\left(\mathbf{x}\right)}\right]\left(\mathbbm{1}-f_{\mathrm{Hxc}}{\chi_{s}}\right)^{-1}\left(\mathbf{x},\mathbf{r}\right)\left.+2\nu_{q}{\tilde{g}_{\mathrm{xc}\,ia,ia}\left[n_{0}\right]\left(\mathbf{x},\omega_{I}\right)}{\chi\left(\mathbf{x},r\right)}\right\}.$
While this expression is formally derived from a single matrix element, it contains information about all KS states of the system through $\chi_{\scriptscriptstyle\rm S}$ and the infinite sums over all KS orbitals $\phi_{p}$. We also define a “single-transition limit” (STL) approximation which keeps only KS orbitals $i$ and $a$ in these terms. STL would be exact in the special case of a two-electron two-level KS system. In this limit, a further simplication ensues: the functions $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}$ and $\chi$ can be exactly defined by the closed-form expressions
$(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}({\bf x},{\bf r})=\delta(\mathbf{x}-{\bf r})-\frac{4}{\nu_{q}+4f_{{\scriptscriptstyle\rm HXC},qq}}f_{\mathrm{HXC},\,q}(\mathbf{x})\Phi_{q}({\bf r})$ (23)
$\chi(\mathbf{x},{\bf r})=\frac{\nu_{q}}{\nu_{q}+4f_{{\scriptscriptstyle\rm HXC},qq}}\chi_{\scriptscriptstyle\rm S}(\mathbf{x},{\bf r}),$ (24)
resulting from a series resummation described in Section 2 of Supplementary Material. Under simply this two-level reduction, one obtains the density difference expression in the following form:
$\Delta n^{\mathrm{STL}}_{I}({\bf r})=$ $\frac{G_{I}^{2}}{\omega_{I}}\left\{\left(\nu_{q}+2f_{{\scriptscriptstyle\rm HXC},qq}(\omega_{I})\right)\Delta n^{\scriptscriptstyle\rm KS}({\bf r})\right.$ (25)
$+\frac{8}{\nu_{q}+4f_{{\scriptscriptstyle\rm HXC},qq}}\left[\left(\nu_{q}+f_{{\scriptscriptstyle\rm HXC},qq}(\omega_{I})\right)\left(f_{\mathrm{HXC},\,ii,\,ia}(\omega_{I})-f_{\mathrm{HXC},\,aa,\,ia}(\omega_{I})\right)\right.$
$-\left.\left.\nu_{q}\tilde{g}_{{\scriptscriptstyle\rm XC},qqq}(\omega_{I})\right]\Phi_{q}({\bf r})\right\},$
We note that no other approximations are applied in derivation of the Eq. (25). For an adiabatic approximation to the kernel, this result is formally equivalent to the density-difference matrix representation given by Ref. 5 in its single-transition limit; see Supplementary Material for the details.
2.1.1 Numerical Evaluation of $\Delta n^{\rm SMA}_{I}({\bf r})$
The calculation of the density difference through Eq. (16) or its SMA approximation through Eq. (22) requires addressing several issues, introducing approximations along the way.
First, we will truncate the infinite sums in Eq. (22) over all KS orbitals ($p$). Since we will be focusing only on the lower excitations, this is not a severe approximation, and convergence can be easily checked.
Second, concerning the $f_{\scriptscriptstyle\rm HXC}$ matrix elements: As noted earlier, the required matrix elements are not only between the pairs of occupied-unoccupied transitions that are the usual ingredients of linear response calculations. While the larger span does not pose a problem for adiabatic $f_{\scriptscriptstyle\rm HXC}$ approximations, computed from the second functional derivative of an energy functional, we note that these matrix elements are largely not known for the existing frequency-dependent $f_{\scriptscriptstyle\rm HXC}$ approximations 21, 27, 23, 24, although we note that the TDDFT kernel derived from the Bethe-Salpeter equation 28, 29, 30 does contain a spatial-dependence that would provide this larger class of matrix elements. Thus when we compute the densities of excited states of double-excitation character, we will instead work directly with the excitation energies provided by the dressed frequency-dependent kernel of Refs. 21, 27, 23, 24, taking their functional derivative directly; we defer a discussion of this to Section 2.2. All other calculations will utilize an adiabatic approximation for both $f_{\scriptscriptstyle\rm HXC}(\omega)$ and $\tilde{g}_{\scriptscriptstyle\rm XC}(\omega)$. This means that the eigenvector squared value $G_{I}^{2}$ in Eq. (22) reduces to $G_{I}^{2}=1$.
The third issue concerns the calculation of the matrix $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}$ in Eq. 22. In non-periodic systems, the values of $\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}$ approach zero at the boundaries of the spatial grid, which serves as a source of instability in many numerical inversion algorithms 31. Further, the direct numerical inversion is computationally expensive, so we will evaluate this by assuming $f_{\scriptscriptstyle\rm HXC}$ is relatively small, and Taylor expand up to first order, in the following way:
$(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}({\bf x},{\bf r})\approx\delta({\bf x}-{\bf r})+\int d^{3}x^{\prime}f_{\scriptscriptstyle\rm HXC}({\bf x},{\bf x}^{\prime})\chi_{\scriptscriptstyle\rm S}({\bf x}^{\prime},{\bf r})$ (26)
$\chi({\bf x},{\bf r})\approx\chi_{\scriptscriptstyle\rm S}({\bf x},{\bf r})+\iint d^{3}x{{}^{\prime}}d^{3}x^{\prime\prime}\chi_{\scriptscriptstyle\rm S}({\bf x},{\bf x}{{}^{\prime}})f_{\scriptscriptstyle\rm HXC}({\bf x}{{}^{\prime}},{\bf x}^{\prime\prime})\chi_{\scriptscriptstyle\rm S}({\bf x}^{\prime\prime},{\bf r})$ (27)
2.2 Densities of States of Double-Excitation Character
With the adiabatic approximation, TDDFT is only able to predict excitations of single excitation character, composed of linear combinations of one electron promoted from an occupied KS state to a virtual one. The formulation for excited-state densities in Refs. 4, 5 is restricted to this case. For states of double-excitation character, a frequency-dependent kernel was developed 20, 21, 22, 23, 24 that has been shown to perform well on a range of systems 32, most recently capturing the curve-crossing between the 1Bu and 2Ag singlet states in trans-butadiene 24, for example. The kernel of Ref. 21, 27 operates within a dressed single-pole approximation (DSPA), generalized to a dressed Tamm-Dancoff approximation (DTDA) when more than one single excitation couples to a double, while, in order to obtain meaningful oscillator strengths as well as energies, the kernel of Ref. 23, 24 operates within a dressed small-matrix approximation (DSMA), generalized to the fully dressed TDDFT (DTDDFT). While Eq. (15) and its SMA Eq (22) hold also beyond the adiabatic approximation, they require matrix elements of the kernel that in a larger space than those derived in the approximations, as discussed above. Therefore, we instead take the functional derivative of Eq. 8 in a different way, directly from the DSPA and DSMA expressions for the excitation energy.
We begin with the DSMA case where 23
$\omega_{I}^{2}{\overset{\rm DSMA}{\approx}}\nu_{q}^{2}+4\nu_{q}f^{\mathrm{DSMA}}_{{\scriptscriptstyle\rm HXC},qq}(\omega),$ (28)
with the frequency-dependent kernel
$f^{\mathrm{DSMA}}_{{\scriptscriptstyle\rm HXC},qq}(\omega)=$ $f^{A}_{{\scriptscriptstyle\rm HXC},qq}+\frac{\left|H_{qD}\right|^{2}}{4\nu_{q}}\left(1+\frac{\left(H_{qq}+H_{DD}-2H_{00}\right)^{2}}{\left[\omega^{2}-\left(\left(H_{DD}-H_{00}\right)^{2}+H_{qD}^{2}\right)\right]}\right),$ (29)
where $f^{A}_{{\scriptscriptstyle\rm HXC},qq}$ is an adiabatic approximation, the matrix elements of the interacting Hamiltonian $H_{mn}$ are taken in the truncated 3-state Hilbert space that contains ground, singly excited $q=i\rightarrow a$ and doubly excited $D=(j\rightarrow b,\,k\rightarrow c)$ states; these states compose the state of interest of double-excitation character, with $\nu_{bj}+\nu_{ck}=({\epsilon}_{b}-{\epsilon}_{j})+({\epsilon}_{c}-{\epsilon}_{k})$ being close to $\nu_{q}={\epsilon}_{a}-{\epsilon}_{i}$, well-separated from other excitations.
Eq. (29) is given for the “0” variant of DSMA Hxc kernel from the work23; other variants reduce the number of two-electron integrals needed by replacing some of the Hamiltonian matrix elements with either Kohn-Sham or appropriate adiabatic TDDFT values. For example, in the “A” variant, $H_{qq}-H_{00}\to\omega_{q}^{\mathrm{ASMA}}$ and $H_{DD}-H_{00}\to\omega_{jb}^{\mathrm{ASMA}}+\omega_{kc}^{\mathrm{ASMA}}$. Here, we have chosen to use the first replacement, but we keep the Hamiltonian matrix element for the double-excitation part. But we note that the variants all give similar results for this system. The two excitation energies given by DSMA Eq. (28) are therefore
$\omega_{\pm}^{2}=\frac{1}{2}\left\{{\omega^{\mathrm{ASMA}}}^{2}+(H_{DD}-H_{00})^{2}+2H_{qd}^{2}\pm\left(\omega^{\mathrm{ASMA}}+H_{DD}-H_{00}\right)\sqrt{\left({\omega^{\mathrm{ASMA}}}-(H_{DD}-H_{00})\right)^{2}+4H_{qD}^{2}}\right\}.$ (30)
Taking the derivative Eq. 8 of Eq.(30) then leads to the DSMA excited state density:
$\Delta n^{\mathrm{DSMA}}_{\pm}({\bf r})$ $=\frac{1}{2}(1\pm\cos\theta)\Delta n^{\mathrm{ASMA}}({\bf r})$ (31)
$+\frac{1}{2\Omega_{\pm}}\left\{\frac{1}{2}(1\mp\cos\theta)\frac{\delta\left(H_{DD}-H_{00}\right)^{2}}{\delta v_{\mathrm{ext}}({\bf r})}\right.$
$\left.+\frac{\delta H_{qD}^{2}}{\delta v_{\mathrm{ext}}({\bf r})}\pm\sin\theta\frac{\delta}{\delta v_{\mathrm{ext}}({\bf r})}\left[H_{qD}\left(\omega^{\mathrm{ASMA}}+(H_{DD}-H_{00})\right)\right]\right\}.$
In the first term above the adiabatic excited state density difference $\Delta n^{\mathrm{ASMA}}$ appears, which we obtain from Eq. (22), with the kernel $f_{\scriptscriptstyle\rm HXC}(\omega_{I})$ approximated by the chosen adiabatic approximation. This density contributes through the mixing angle $\theta$ defined via
$\tan\theta=\frac{2H_{qD}}{\omega^{\mathrm{ASMA}}-(H_{DD}-H_{00})}$ (32)
that scales linearly with the coupling between states $q$ and $D$.
Note that in the first term of Eq. (31), the adiabatic density-difference $\Delta n^{\mathrm{ASMA}}({\bf r})$ is scaled by the square of the Casida eigenvector $G_{\pm}^{2}$ from Eq. (21),
$G_{\pm}^{2}=\frac{1}{2}\left(1\pm\cos\theta\right)\,,$ (33)
which is a measure of the amount of single-excitation character of the state 23. In the limit $\cos\theta=1$, the state indexed by $+$ reduces to a pure double excitation while that indexed by -- reduces to a pure single excitation; the opposite is true for $\cos\theta=-1$. On the other hand, when $\cos\theta\to 0$, the states gain an equal contribution of single and double excitation character, and become nearly degenerate. Their densities however remain distinct, due to the last term in Eq. (31).
For the second and third terms of Eq. (31), which contain non-adiabatic contributions to the density difference, one needs derivatives of the Hamiltonian matrix elements. These can be evaluated from evaluating the Hamiltonian matrix elements via Slater-Condon rules for one-body and two-body integrals and then taking the derivatives (details in Supplementary Material):
$\frac{\delta}{\delta v_{\rm ext}({\bf r})}\langle\phi_{r}|\hat{h}|\phi_{s}\rangle$ $=\Phi_{rs}({\bf r})$ (34)
$+\int d^{3}x\left\{\sum_{p\neq r}^{\infty}\frac{\langle\phi_{p}|\hat{h}|\phi_{s}\rangle}{{\epsilon}_{r}-{\epsilon}_{p}}\Phi_{pr}({\bf x})+\sum_{p\neq s}^{\infty}\frac{\langle\phi_{r}|\hat{h}|\phi_{p}\rangle}{{\epsilon}_{s}-{\epsilon}_{p}}\Phi_{ps}({\bf x})\right\}\left(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}\right)^{-1}({\bf x},{\bf r}),$
$\frac{\delta}{\delta v_{\rm ext}({\bf r})}(\phi_{r}\phi_{s}|\phi_{m}\phi_{n})$ $=\int d^{3}x\left\{\sum_{p\neq r}^{\infty}\frac{(\phi_{p}\phi_{s}|\phi_{m}\phi_{n})}{{\epsilon}_{r}-{\epsilon}_{p}}\Phi_{pr}({\bf x})+\sum_{p\neq s}^{\infty}\frac{(\phi_{r}\phi_{p}|\phi_{m}\phi_{n})}{{\epsilon}_{s}-{\epsilon}_{p}}\Phi_{ps}({\bf x})\right.$ (35)
$+\left.\sum_{p\neq m}^{\infty}\frac{(\phi_{r}\phi_{s}|\phi_{p}\phi_{n})}{{\epsilon}_{m}-{\epsilon}_{p}}\Phi_{pm}({\bf x})+\sum_{p\neq n}^{\infty}\frac{(\phi_{r}\phi_{s}|\phi_{m}\phi_{p})}{{\epsilon}_{n}-{\epsilon}_{p}}\Phi_{pn}({\bf x})\right\}\left(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}\right)^{-1}({\bf x},{\bf r}),$
where $\hat{h}=-\frac{1}{2}\nabla^{2}+\hat{v}_{\rm ext}$, and the notations are chosen so that $\langle\phi_{r}|\hat{h}|\phi_{s}\rangle=\int d^{3}x\phi_{r}({\bf x})h({\bf x})\phi_{s}({\bf x})$ and $(\phi_{r}\phi_{s}|\phi_{m}\phi_{n})=\iint d^{3}xd^{3}x{{}^{\prime}}\phi_{r}({\bf x})\phi_{s}({\bf x}{{}^{\prime}})w({\bf x},{\bf x}{{}^{\prime}})\phi_{m}({\bf x}{{}^{\prime}})\phi_{n}({\bf x}{{}^{\prime}})$.
While including both excitation and de-excitation components to the TDDFT matrix is necessary to obtain meaningful transition densities 23, we will check whether including only the excitation components is adequate for the excited state density itself, by using the Tamm-Dancoff approximation of DSMA. This is just the dressed single-pole approximation (DSPA) that was derived earlier in Ref. 21:
$\omega_{I}\overset{DSPA}{\approx}\nu_{q}+2f_{{\scriptscriptstyle\rm HXC},qq}^{\mathrm{DSPA}}(\omega_{I}),$ (36)
where the DSPA kernel is
$f_{{\scriptscriptstyle\rm HXC},qq}^{\mathrm{DSPA}}(\omega)=f_{{\scriptscriptstyle\rm HXC},qq}^{A}+\frac{1}{2}\frac{\left|H_{qD}\right|^{2}}{\omega-\left(H_{DD}-H_{00}\right)}.$ (37)
The excited-state density difference following from Eq. (36) is
$\Delta n^{\mathrm{DSPA}}_{\pm}({\bf r})$ $=\frac{1}{2}\left(1\pm\cos\theta\right)\Delta n^{\mathrm{ASPA}}({\bf r})$ (38)
$+\frac{1}{2}\left(1\mp\cos\theta\right)\frac{\delta(H_{DD}-H_{00})}{\delta v_{\rm ext}({\bf r})}$
±$\pm\sin\theta\frac{\delta H_{qD}}{\delta v_{\rm ext}({\bf r})},$
where $\Delta n^{\mathrm{ASPA}}({\bf r})$ is also derived from the derivative of the excited-state energy within the adiabatic SPA:
$\Delta n_{I}^{\mathrm{ASPA}}(\mathbf{r})$ $=\int d^{3}x\left\{\left[\Delta n_{I}^{\mathrm{KS}}(\mathbf{x})\right.+4\sum_{p\neq a}^{\infty}\frac{1}{\epsilon_{a}-\epsilon_{p}}{f_{\mathrm{Hxc}\,ip,ia}}{\Phi_{pa}\left(\mathbf{x}\right)}\right.$ (39)
$\left.-4\sum_{p\neq i}^{\infty}\frac{1}{\epsilon_{p}-\epsilon_{i}}{f_{\mathrm{Hxc}\,pa,ia}}{\Phi_{ip}\left(\mathbf{x}\right)}\right]\left(\mathbbm{1}-f_{\mathrm{Hxc}}{\chi_{s}}\right)^{-1}\left(\mathbf{x},\mathbf{r}\right)\left.+2{g_{\mathrm{xc}\,ia,ia}\left[n_{0}\right]\left(\mathbf{x}\right)}{\chi\left(\mathbf{x},r\right)}\right\}.$
3 Results and Analysis on Model Systems
We now apply our method to a set of one-dimensional models of two soft-Coulomb interacting electrons with different external potentials. We define the three following external potentials: (i) Helium $v_{\rm ext}^{\mathrm{1DHe}}(x)=\frac{-2}{\sqrt{1+x^{2}}}$ to model local excitations, (ii) two double-well models, one in which the ground-state density is largely localized in a soft-Coulomb potential on the left, $v_{\rm ext}^{\mathrm{gs-soft}}(x)=-\frac{2}{\sqrt{(x+R/2)^{2}+1}}-\frac{1}{\cosh^{2}(x-R/2)}$ 33, and the other in which the ground-state density is more tightly localized in the left well, $v_{\rm ext}^{\mathrm{gs-loc}}(x)=-\frac{2}{\sqrt{(x+R/2)^{2}+1}}-\frac{2.9}{\cosh^{2}(x+R/2)}-\frac{1}{\cosh^{2}(x-R/2)}$, where $R=7$a.u., and in these wells we will model charge-transfer excitations, and (iii) perturbed harmonic $v_{\rm ext}^{\mathrm{Harm_{\gamma}}}(x)=\frac{1}{2}x^{2}+\gamma|x|$ for modeling the double excitations, where $\gamma$ parameter tunes the degree of mixing between a single excitation and double-excitation in the second multiplet.
With just two electrons, the Kohn-Sham ground-state has both electrons occupying the lowest-energy orbital $i=1$, and the KS response function for all three cases can be written as
$\chi_{\scriptscriptstyle\rm S}(x,x{{}^{\prime}})=-4\sum_{a}^{\mathrm{unocc}}\frac{\phi_{i}(x)\phi_{a}(x)\phi_{a}(x{{}^{\prime}})\phi_{i}(x{{}^{\prime}})}{\epsilon_{a}-\epsilon_{i}}.$ (40)
The simplicity of the models enables the computations of the exact KS orbitals, and we can compare the effect of using approximate versus exact KS orbitals in the expressions for the excited-state densities. We will first consider the lowest energy excitations in 1DHe and the two double-wells. These are dominated by single excitations, and so we will consider two adiabatic approximations: exact-exchange (EXX) and local-density approximation (LDA) 34, which are first-principles approximations at two opposite extremes, the former derived from many-body perturbation theory has non-local density-dependence and the latter derived from the uniform electron gas paradigm has completely local density-dependence. We will compare the fully-converged SMA densities (Eq. 22) with those obtained from STL (Eq. 25). For the Harmγ system, we compare the performance of ASMA using the EXX approximation with the DSMA Eq. (31).
For our calculations of the exact ground and excited-state densities, and the exact, Kohn-Sham, EXX, and LDA orbitals of these one-dimensional systems, we use the Octopus code35. The exact Kohn-Sham potential and orbitals are obtained from first inverting the ground-state Kohn-Sham equation using the exact ground-state density of a two-electron system, and then finding its eigenstates and eigenvalues in an in-house code. The helium model was simulated in a box ranging from -40 to 40 a.u. with a grid spacing of 0.1 a.u. For both double-well models, a box from -50 to 50 a.u. with the same grid spacing was used. Finally, for the system in the harmonic potential, simulations were performed in a box from -20 to 20 a.u. with a grid spacing of 0.05 a.u.
3.1 One-Dimensional Helium: Local Excitations
Figure 1 shows the excited-state density differences of the first four local excitations of 1DHe. In this plot, the exact KS orbitals are used in Eq. 22 and $f_{\scriptscriptstyle\rm HXC}$ and $\tilde{g}_{\scriptscriptstyle\rm XC}$ are approximated by EXX: $f_{\scriptscriptstyle\rm HXC}(x,x^{\prime})=f_{\scriptscriptstyle\rm HX}(x,x^{\prime})=\frac{1}{2\sqrt{1+(x-x^{\prime})^{2}}}$, and $\tilde{g}_{\scriptscriptstyle\rm XC}=\tilde{g}_{\scriptscriptstyle\rm X}=0$. The plots show the fully converged SMA and the STL excited-state density differences, as well as the KS one; the inset shows the excited-state density itself, obtained from adding the exact ground-state density to the density-difference in each case.
Figure 1: Excited-state density differences for 1D Helium: TDDFT within SMA, STL, compared with exact and KS, for the lowest 4 excitations (panels (a)–(d)). The insets show the exact excited-state density in each case. For the approximations, the exact KS orbitals are used in the formulas for the densities, and $f_{\scriptscriptstyle\rm XC}^{\rm EXX}$ is used for the kernel. Panel (e) shows the convergence for the lowest 3 excited densities with respect to the number of orbitals included in the sums over $p$.
We see that SMA gives a noticeable correction to the KS excited-state densities, particularly for the lowest excitation, and that the SMA, STL, and KS densities all become more accurate for the higher excitations, as expected. In particular, for the lowest excitation, SMA captures the sharper peak and the shoulders around the location of the He nucleus, which are missed by the KS density in the lowest excitation, as seen in the inset. For this lowest excitation, the effect of the higher KS orbitals is important, since the STL density is almost indistinguishable from the KS density. For the higher excitations shown, STL brings a small correction to the KS density (in fact, almost none for the third excited state).
[…]
In Figure 2, we compare the SMA densities predicted with the EXX and LDA approximations for the orbitals and the $f_{\scriptscriptstyle\rm HXC}$ kernel; we take again $\tilde{g}_{\scriptscriptstyle\rm XC}=\tilde{g}_{\scriptscriptstyle\rm X}=0$ throughout. We find that EXX performs notably better than LDA, particularly for the choice of the KS orbitals. Using $f_{\scriptscriptstyle\rm HXC}^{\rm LDA}$ with EXX orbitals tends to slightly under- and over-estimate the density difference near the position of He nucleus, but the effect of the kernel is less than the effect of the choice of the KS orbital approximation. Though SMA with LDA orbitals seems to estimate the density difference only slightly worse for the lowest excitation, for higher ones, it fails to capture the density difference accurately away from the nucleus, as might be expected since the too-rapid decay of the LDA KS potential causes significant errors in the KS orbitals away from the core.
Figure 2: Excited-state density differences for 1D Helium: comparison of approximations for KS orbitals and $f_{\scriptscriptstyle\rm HXC}$ within SMA for the lowest 4 excitations.
3.2 Double-well potentials: Charge-Transfer Excitations
We next turn to the lowest charge-transfer excitations in the double-well potential systems. The parameters chosen are such that the ground-state density is localized in the left well.
We begin with the double-well where the ground-state has both electrons in the soft-Coulomb potential on the left ($v_{\rm ext}^{\mathrm{gs-soft}}$), that had been studied in Ref. 33. The lowest excitation has a charge-transfer nature. In Fig. 3(a), we observe that, when exact KS orbitals are used to evaluate the sums in the expressions, the KS density is already a good approximation with only a small error to the density of the charge-transfer excitation, that STL gives practically no correction, while the SMA gives a larger correction, giving a density very close to the exact. The results in this figure used EXX for the kernel.
Figure 3: Excited-state density differences for the CT excitation of FERM13 double well model33. Panel (a) shows SMA, STL compared to the exact and KS excited-state density difference. Exact KS orbitals and $f_{\scriptscriptstyle\rm XC}^{\mathrm{EXX}}$ we used in approximations. The inset shows the excited state density of CT state. Panel (b) shows the performance of SMA with KS orbitals and $f_{\scriptscriptstyle\rm XC}$ approximated by LDA and EXX functionals. Panel (c) illustrates the issues of SMA and STL excited-state density differences when . Panel (d) shows the exact, EXX ground state densities, along with and LDA ground state which is wrongfully localized in the second well, and the corresponding KS potentials in the inset.
In Fig. 3(b), we compare the performance of approximations for the KS orbitals and the kernel. An immediate observation is that the use of LDA orbitals leads to huge errors with any choice of kernel approximation. Using EXX orbitals gives good results similar to using exact KS orbitals, when either the EXX or LDA kernels are used, with the former kernel being slightly more accurate. In contrast, the use of LDA orbitals leads to completely wrong density-differences for the charge-transfer excitation, even yielding negative density in some regions of space, and divergences at the two nuclei.
Why does this problem emerge? LDA underestimates the gap between the occupied and lowest unoccupied KS levels: while the exact LUMO-HOMO gap is $\nu_{\mathrm{exact}}=0.112$ a.u., the LDA gap is far smaller, at $\nu_{\mathrm{LDA}}=0.005$ a.u. This means that for all excitations, the terms in Eq. (22) have much too small denominators, which in itself would lead to divergences unless cancelled out by a similar term in the numerator. For this lowest charge-transfer excitation, this cancellation does in fact occur, but the divergence remains because the gap appears also in the denominator of $\chi_{\scriptscriptstyle\rm S}$. When we approximate $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}$ by its first-order term in the Taylor expansion in $f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}$, the severely underestimated denominator leads to the divergence. Thus, the problem lies not with the SMA itself, but rather with our approximation of the inverse of $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}$ that appears in it.
Instead, as discussed earlier, the STL approach resums the Taylor series (Eq. (23)) and so treats this inverse exactly, although within the STL approximation. In Fig. 3(c), we show that the STL density difference with LDA indeed does not display any divergence, while approximating the inverse function $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}$ by its first-order term in $f_{\scriptscriptstyle\rm HXC}$ in the same way as done in SMA does diverge. However, with LDA, the STL charge-transfer density-difference still gives substantial error compared to exact one, and this error stems from the erroneous LDA ground-state shown in Fig. 3(d). The LDA suffers from a delocalization error, and has an unphysical fractional charge in each well instead of asymptotically having a two-electron density fully on the left. The ground-state fractional charge error comes hand in hand with the vanishing LDA KS HOMO-LUMO gap, and this fractional-charge error is transferred to the KS charge-transfer excitation too. This is evident in Fig. 3(c) where integrating the KS curve on either side of zero falls short of 1. The wrong behavior of the KS orbitals leads to an erroneous STL density difference, even with the wrong sign.
To emphasize the importance of an accurate KS ground-state, we perform the same calculations on the double-well model with the potential $v_{\rm ext}^{\mathrm{gs-loc}}$ from Ref. 36, where the two electrons in the ground-state are more tightly localized in the left well, so that LDA has a better hope to capture this qualitatively correctly. The second lowest excitation of this system is the charge transfer, the density of which is shown in Fig. 4.
Figure 4: Excited-state density differences for the CT excitation in the double well model $v_{\rm ext}^{\mathrm{gs-loc}}$. Panel (a) shows SMA, STL compared to the exact and KS excited-state density difference, using the exact KS orbitals and $f_{\scriptscriptstyle\rm XC}^{\mathrm{EXX}}$. The inset shows the excited state density of CT state. Panel (b) shows the performance of SMA with KS orbitals and $f_{\scriptscriptstyle\rm XC}$ approximated by LDA and EXX functionals as indicated in the legend. Panel (c) illustrates the performance of STL compared to SMA and exact excited-stated density difference. Panel (d) shows the exact, EXX and LDA ground state densities, and corresponding KS potentials in the inset.
The charge-transfer excited state density-differences obtained with exact KS orbitals and $f_{\scriptscriptstyle\rm HXC}^{\mathrm{EXX}}$ (Fig. 4(a)), as well as with EXX and LDA approximations (Fig. 4(b)), are all in good agreement with the exact. We highlight specifically the result given by LDA approximation, since unlike in the previous result, we not only do not encounter the divergences of the density difference, but it even agrees closely with the exact result. The KS excitation frequency is not qualitatively different from the exact, as it was in the previous case: the exact HOMO-LUMO gap is $\nu_{\mathrm{exact}}=2.235$ a.u., while LDA one is $\nu_{\mathrm{LDA}}=1.974$ a.u. We emphasize that getting reasonable KS quantities is of the most importance for the good performance of the excited state densities. Regarding the approximate evaluations of these densities, we observe that while both SMA and STL accurately capture the charge transfer in this case, STL performs slightly better. This can be explained by inspecting the behavior of LDA KS potential in the inset of Fig. 4(d): the lowest KS orbitals tend to be more accurate, leading to a good approximation of the ground state, but a worse approximation of excited-state density difference (second-excited state) with SMA.
3.3 Harmonic potentials: Double excitations
Turning now to our third model, Harmγ, our final example is that of excited-state densities of states of double-excitation character. In this system the lowest excitation is a single excitation, and we expect adiabatic SMA to work particularly well especially since it is well-separated from other excitations. The 2nd and 3rd excitations are mixtures of KS single and double excitations; at $\gamma=0$ the mixture is close to 50:50, so that the first and the second terms of Eq. (31) and Eq. (38) contribute equally to the density-difference of both states while as $\gamma$ increases to one, the KS near-degeneracy is increasingly lifted, with the lower state of the pair acquiring a predominantly single excitation character, and the upper state a predominantly double.
Fig. 5 shows the excited-state densities of the lowest three excitations, predicted from adiabatic SMA, SPA, frequency-dependent DSMA and DSPA, and compared with the exact. As expected, the approximations coincide closely with the exact for the lowest excitation, but differ for the second and third due to their double-excitation character. Both DSPA and DSMA do an excellent job, while the adiabatic SMA and SPA qualitatively fail for both states at $\gamma=0$. For the third state at $\gamma=1$ the adiabatic approximations also fail, but they provide a closer approximation for the 2nd state, with a small error at the origin, consistent with the fact that it is dominated by the single excitation. We note there is only one adiabatic state, so the predicted density for both the actual 2nd and 3rd states is the same with the adiabatic SMA/SPA.
Figure 5: Column 1 - 1st excitation of 1d Harmonic oscillator at degrees of mixing $\gamma=0$ and $\gamma=1$, Exact vs. SMA; Column 2 - 1st of the double excitations at $\gamma=0$ and $\gamma=1$, DSMS vs SMA vs Exact; Column 3 - 2nd of the double excitations at $\gamma=0$ and $\gamma=1$, DSMS vs SMA vs Exact.
Thus, we have shown that the frequency-dependent kernels of Ref. 21, 23, which were shown earlier to produce good excitation energies for states of double-excitation character, and that of Ref. 23 also good oscillator strengths, both also yield excellent approximations to excited-state densities.
4 Conclusions
In this work we developed a general formalism for obtaining the real-space excited-state densities from linear response TDDFT using time-independent perturbation theory. While the approach is equivalent to earlier work 4, 5 presented using the stationary property of excitation energies and the variational principle, our work here goes beyond the restriction to the adiabatic approximation assumed there. We observe that even with two contrasting approximations (LDA and EXX), we obtain accurate results for the lowest local and charge-transfer excitations between closed-shell fragments. Within the small-matrix approximation, which in its essence retains a single KS transition in its calculation of the energy, the excited-state density still incorporates corrections from all possible KS transitions, leading to a significant improvement over the KS approximation. Our results show the importance of an accurate KS potential in the calculation of excited-state density, which is better provided by adiabatic EXX than the LDA. For the future calculations of excited-state densities in real systems, our results are consistent with previous works 8, 9, 10, 11, 12, 13, 14 that suggest the use of hybrid functionals with large fraction of exact exchange as a more reliable choice.
The practical use of the general formalism is limited to the adiabatic approximation and to proposed approximations that are available for frequency-dependent $f_{\scriptscriptstyle\rm HXC}$ kernels. (A general-purpose frequency-dependent kernel does not exist). We show that for doubly-excited states, the densities from the DSPA and DSMA-based formulation, both previously shown to be good for excitation energies and DSMA for oscillator strengths, provides accurate double excitation densities, which is not achievable through adiabatic approximation.
Benchmarking on a wider set of real molecules is required to further probe the efficiency and accuracy of the linear-response formalism developed for real-space excited-state densities for states of both single and double excitation character, as has been explored for dipole moments restricted to the adiabatic approximation (single excitations) in earlier works 8, 9, 10, 11, 12, 13, 14, 15. Beyond static properties, the formalism can be used in the response-reformulated TDDFT for nonequilibrium dynamics 37, a key ingredient of which are the densities of the states involved in the dynamics. In this context, it is of particular interest to explore its performance in fully non-equilibrium dynamics which have a propensity to occupy states of double-excitation character during the evolution, whose densities can be obtained from our formulation.
{acknowledgement}
Financial support from the National Science Foundation Award CHE-2154829 (AB) and the Department of Energy, Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences and Biosciences under Award No. DE‐SC0024496 (NTM) is gratefully acknowledged.